# Cross-Modal Bidirectional Interaction Model for Referring Remote Sensing Image Segmentation

### v1

@@ V-C3 RisBench
Compared to other models utilizing the Swin-B and BERT combinations, CroBIM significantly enhances cross-modal alignment between visual and textual features through improvements in feature extraction. By leveraging an innovative cross-modal bidirectional interaction mechanism, our model more effectively captures key information in the image that corresponds to the textual description. This capability is particularly beneficial in remote sensing imagery where targets are complex and the background is often noisy. Even under such challenging conditions, CroBIM consistently achieves excellent segmentation results. The visualization results in Fig. 11 further highlight the superiority of our approach.
VI Ablation Study
@@
VI-A Design of CAPM
To further validate the effectiveness of the CAPM module, we conduct a detailed analysis of its context-aware prompt design. First, we performe ablation studies on the combinations of visual features at different scales, with the results presented in Table. V(a). As can be observed, the combination incorporating all four visual features, ${V_{1},V_{2},V_{3},V_{4}}$, achieves the best performance in terms of prediction accuracy (Pr@0.5, Pr@0.7, Pr@0.9) and mIoU. This demonstrates that integrating visual features from multiple scales significantly enhances the model’s prediction accuracy and overall segmentation quality. The improved performance suggests that the comprehensive utilization of features across different levels of granularity contributes to a more robust and precise representation, thus improving the model’s ability to capture complex visual cues.
We further investigate the impact of different pooling strategies in the dimensionality reduction of multi-scale visual features, as shown in Table. V(b). It can be observed that, compared to max pooling and average pooling, adaptive average pooling consistently achieves superior performance across all evaluation metrics. These results suggest that the adaptive mechanism is more effective in flexibly integrating features across different scales, thereby preserving more information that is crucial for accurate predictions. Moreover, by dynamically adjusting the weighting of features based on their respective scales, adaptive average pooling enables a more comprehensive retention of both global and local information during the dimensionality reduction process, ultimately enhancing segmentation performance.
TABLE V: Ablation studies on options design of CAPM.
@@ VI-A Design of CAPM
TABLE VI: Ablation studies on attention deficit compensation of LGFA.
@@
VI-B Design of LGFA
We further validated the effectiveness of attention deficit compensation within the LGFA module. The ablation study results presented in Table VI demonstrate the efficacy of applying attention deficit compensation at different stages (S1-S4) in improving the aggregation and alignment of visual and linguistic features. As attention deficit compensation is progressively applied from a single stage (S4) to multiple stages, a significant performance improvement across all metrics can be observed. Notably, when attention deficit compensation is applied at all stages (S1-S4), the model achieves the best segmentation results: Pr@0.5 of 77.55%, Pr@0.7 of 66.38%, and mIoU of 69.33%. The most prominent gains are observed in Pr@0.5 and oIoU, indicating that the complete attention deficit compensation strategy enhances the precision of feature alignment and overall segmentation accuracy. These findings underscore the importance of applying attention deficit compensation at multiple stages, as it leads to more robust integration of visual and linguistic features, thereby improving the segmentation performance.
VI-C Design of MID
Moreover, we investigat the effectiveness of the cascaded bidirectional attention mechanism (CBAM) employed in our MID, and compared it with two existing approaches: the unidirectional visual-to-language attention module PWAM[19], and the parallel bidirectional attention WPA[57]. The comparative results are presented in Table. VII. As observed, the unidirectional attention mechanism in PWAM yielded the lowest segmentation performance. This underperformance can be attributed to its reliance on unidirectional interactions, which fail to adequately capture the joint representation of cross-modal visual-linguistic features.
In contrast, our cascaded bidirectional attention mechanism achieved the best performance, outperforming PWAM and WPA by 3.79% and 2.31% in terms of mIoU, respectively. These results demonstrate that the cascaded bidirectional attention effectively enables deep interaction between linguistic features and multi-scale visual contexts, fostering cross-modal alignment in a more efficient manner. This enhanced interaction not only improves the quality of feature representation but also significantly boosts segmentation performance, highlighting the superiority of our proposed approach in addressing the challenges of visual-linguistic cross-modal tasks.
TABLE VII: Ablation studies on attention mechanisms of MID.
@@ VI-C Design of MID
VI-D Qualitative Results
@@
In Fig. 12, we visualize the attention maps at various stages of the model to analyze CroBIM’s cross-modal alignment mechanism in greater depth. First, in the encoder’s attention maps, we observe that as the network depth increases, CroBIM progressively focuses on more fine-grained target areas. The attention maps from the early stages reveal a broad feature capture, with attention distributed across the global information in the image. This is closely related to the high resolution and complex background of remote sensing imagery—at these early stages, the model must focus on the overall structure to establish a comprehensive understanding of the scene. As the network layers progress, attention increasingly narrows to target regions that are closely aligned with the textual description, indicating that the model effectively filters out background noise and captures the key features of the target. For instance, in remote sensing images with complex terrain or buildings, CroBIM is able to accurately localize critical objects such as buildings and roads in the mid-to-late layers. Furthermore, the attention maps in the decoder stage illustrate the model’s segmentation capability. In the decoder, the attention maps demonstrate a high degree of refinement, with the model exhibiting more precise attention to the edges and finer details of the target regions. This focus further validates CroBIM’s strong performance in processing high-resolution remote sensing images, particularly in tasks requiring precise target segmentation. The decoder’s attention maps clearly show that the model can accurately delineate the boundaries between the target and background, especially in areas where textures or colors are highly similar.
Figure 13: Failure cases of our proposed CroBIM on the RisBench test set.
@@
Although our CroBIM effectively models the query expression and aligns visual features with textual embeddings, some failure cases still occur, as shown in Fig. 13. These failure modes can be categorized into four types. First, as illustrated in the first column, when the textual description refers to multiple objects within the image, semantic ambiguity can lead to the model segmenting multiple objects instead of the intended one. Second, as seen in the second column, when the foreground object and background share similar visual appearances, the model is prone to errors. The third column presents another common failure type—imprecise annotations. Lastly, when the target is located near the image boundary, as shown in the fourth column, the lack of sufficient contextual information may cause the model to misidentify the target’s location and boundaries, leading to inaccurate segmentation results.
VII Conclusion
@@
For future work, we aim to embed domain-specific knowledge[58] of remote sensing imagery into language models, such as sensor imaging theory, spatial correlations, and spectral characteristics of ground objects, to further enhance remote sensing data analysis and interpretation. Additionally, another promising research direction is the integration of textual information with remote sensing through geolocation. By incorporating non-traditional geographic data, such as geotagged social media posts[59] and newspaper articles, remote sensing data can be combined with complementary sources. This approach broadens the potential applications[60, 61, 62] of remote sensing visual-language foundation models.

### v2

@@ 5.3.3 RisBench
Compared to other models utilizing the Swin-B and BERT combinations, CroBIM significantly enhances vision-language alignment between visual and textual features through improvements in feature extraction. By leveraging an innovative cross-modal bidirectional interaction mechanism, our model more effectively captures key information in the image that corresponds to the textual description. This capability is particularly beneficial in remote sensing imagery where targets are complex and the background is often noisy. Even under such challenging conditions, CroBIM consistently achieves excellent segmentation results. The visualization results in Fig. 11 further highlight the superiority of our approach.
Besides, the experimental results demonstrated that our proposed CroBIM framework not only outperformed the existing RRSIS-specific methods (LGCE, RMSIN and MAFN) but also surpassed the adapted SOTA methods from computer vision. This highlighted the effectiveness and robustness of our approach in handling the unique challenges of remote sensing imagery, such as complex spatial relationships and diverse object scales.
6 Ablation Study
@@
6.1 Effectiveness of Each Components
To evaluate the effectiveness of the three key components of CroBIM—CAPM, LGFA, and MID—we conducted ablation experiments on the RISBench test dataset. The experimental results are presented in Fig. 12.
When the CAPM module is removed from the complete CroBIM framework, the mIoU drops from 69.33% to 65.57%, representing a significant performance reduction of 3.76%. Similarly, the oIoU decreases from 73.04% to 71.88%. These results underscore the critical role of CAPM in integrating multi-scale spatial context into linguistic features, thereby enabling precise localization of target objects in remote sensing images. In the absence of CAPM, the model struggles to capture the spatial positional relationships described in the referring expressions, leading to a noticeable decline in segmentation accuracy.
Figure 12: Performance comparison of CroBIM and its ablated variants (w/o CAPM, LGFA, and MID) on the RISBench test set.
When the LGFA module is further removed, the performance degradation becomes more pronounced. The mIoU decreases to 61.28% (a reduction of 4.29%) and the oIoU drops to 70.95% (a reduction of 0.93%) compared to the prior configuration. This highlights the pivotal role of LGFA in aggregating visual and linguistic features across multiple scales and compensating for attention deficits. Without LGFA, the model loses its ability to effectively handle objects of varying scales, which is a fundamental challenge in remote sensing scenarios.
Finally, replacing the MID module with a simple decoder[51] results in mIoU and oIoU dropping further to 54.76% and 68.21%, respectively. These findings clearly demonstrate the importance of MID in ensuring precise alignment between visual and linguistic features through iterative bidirectional refinement. The absence of MID significantly impairs the model’s ability to resolve the complex cross-modal relationships required for accurate segmentation, particularly in remote sensing datasets characterized by diverse geospatial relationships and linguistically rich referring expressions.
In summary, the above results conclusively demonstrate that all three components are indispensable for achieving precise segmentation in complex scenarios. Each module contributes uniquely to the overall performance of CroBIM, ensuring its robustness and effectiveness in addressing the challenges inherent in RRSIS tasks.
Table 5: Ablation studies on options design of CAPM.
@@ 6.1 Effectiveness of Each Components
6.2 Design of CAPM
To further validate the effectiveness of the CAPM module, we conduct a detailed analysis of its context-aware prompt design. First, we performe ablation studies on the combinations of visual features at different scales, with the results presented in Table. 5(a). As can be observed, the combination incorporating all four visual features, ${V_{1},V_{2},V_{3},V_{4}}$, achieves the best performance in terms of prediction accuracy (Pr@0.5, Pr@0.7, Pr@0.9) and mIoU. This demonstrates that integrating visual features from multiple scales significantly enhances the model’s prediction accuracy and overall segmentation quality. The improved performance suggests that the comprehensive utilization of features across different levels of granularity contributes to a more robust and precise representation, thus improving the model’s ability to capture complex visual cues.
We further investigate the impact of different pooling strategies in the dimensionality reduction of multi-scale visual features, as shown in Table. 5(b). It can be observed that, compared to max pooling and average pooling, adaptive average pooling consistently achieves superior performance across all evaluation metrics. These results suggest that the adaptive mechanism is more effective in flexibly integrating features across different scales, thereby preserving more information that is crucial for accurate predictions. Moreover, by dynamically adjusting the weighting of features based on their respective scales, adaptive average pooling enables a more comprehensive retention of both global and local information during the dimensionality reduction process, ultimately enhancing segmentation performance.
6.3 Design of LGFA
We further validated the effectiveness of attention deficit compensation within the LGFA module. The ablation study results presented in Table 6 demonstrate the efficacy of applying attention deficit compensation at different stages (S1-S4) in improving the aggregation and alignment of visual and linguistic features. As attention deficit compensation is progressively applied from a single stage (S4) to multiple stages, a significant performance improvement across all metrics can be observed. Notably, when attention deficit compensation is applied at all stages (S1-S4), the model achieves the best segmentation results: Pr@0.5 of 77.55%, Pr@0.7 of 66.38%, and mIoU of 69.33%. The most prominent gains are observed in Pr@0.5 and oIoU, indicating that the complete attention deficit compensation strategy enhances the precision of feature alignment and overall segmentation accuracy. These findings underscore the importance of applying attention deficit compensation at multiple stages, as it leads to more robust integration of visual and linguistic features, thereby improving the segmentation performance.
Table 6: Ablation studies on attention deficit compensation of LGFA.
@@ 6.3 Design of LGFA
Table 7: Ablation studies on attention mechanisms of MID.
@@
6.4 Design of MID
Moreover, we investigat the effectiveness of the cascaded bidirectional attention mechanism (CBAM) employed in our MID, and compared it with two existing approaches: the unidirectional visual-to-language attention module PWAM[51], and the parallel bidirectional attention WPA[60]. The comparative results are presented in Table. 7. As observed, the unidirectional attention mechanism in PWAM yielded the lowest segmentation performance. This underperformance can be attributed to its reliance on unidirectional interactions, which fail to adequately capture the joint representation of cross-modal visual-linguistic features.
In contrast, our cascaded bidirectional attention mechanism achieved the best performance, outperforming PWAM and WPA by 3.79% and 2.31% in terms of mIoU, respectively. These results demonstrate that the cascaded bidirectional attention effectively enables deep interaction between linguistic features and multi-scale visual contexts, fostering vision-language alignment in a more efficient manner. This enhanced interaction not only improves the quality of feature representation but also significantly boosts segmentation performance, highlighting the superiority of our proposed approach in addressing the challenges of visual-linguistic cross-modal tasks.
Figure 14: Comparison of attention maps between CroBIM and RMSIN at the decoder stage.
6.5 Qualitative Results
@@
In Fig. 13, we visualize the attention maps at various stages of the model to analyze CroBIM’s vision-language alignment mechanism in greater depth. First, in the encoder’s attention maps, we observe that as the network depth increases, CroBIM progressively focuses on more fine-grained target areas. The attention maps from the early stages reveal a broad feature capture, with attention distributed across the global information in the image. This is closely related to the high resolution and complex background of remote sensing imagery—at these early stages, the model must focus on the overall structure to establish a comprehensive understanding of the scene. As the network layers progress, attention increasingly narrows to target regions that are closely aligned with the textual description, indicating that the model effectively filters out background noise and captures the key features of the target. For instance, in remote sensing images with complex terrain or buildings, CroBIM is able to accurately localize critical objects such as buildings and roads in the mid-to-late layers. Furthermore, the attention maps in the decoder stage illustrate the model’s segmentation capability. In the decoder, the attention maps demonstrate a high degree of refinement, with the model exhibiting more precise attention to the edges and finer details of the target regions. This focus further validates CroBIM’s strong performance in processing high-resolution remote sensing images, particularly in tasks requiring precise target segmentation. The decoder’s attention maps clearly show that the model can accurately delineate the boundaries between the target and background, especially in areas where textures or colors are highly similar.
Figure 15: Failure cases of our proposed CroBIM on the RisBench test set. From left to right, the four failure types are semantic ambiguity, visual similarity, imprecise annotations, and boundary targets.
To further demonstrate the superiority of CroBIM, we compared its attention maps with those of RMSIN at the final stage of the decoder, as shown in Fig. 14. The comparison reveals that RMSIN often suffers from attention drift, failing to focus accurately on the target regions specified in the textual query. In contrast, CroBIM consistently maintains precise attention on the target objects, even in complex scenarios with challenging geospatial relationships or similar foreground-background appearances. This refined attention mechanism enables CroBIM to achieve high-precision segmentation by accurately delineating target boundaries and capturing fine details, particularly in regions with similar textures or colors.
@@
Although our CroBIM effectively models the query expression and aligns visual features with textual embeddings, some failure cases still occur, as shown in Fig. 15. These failure cases can be categorized into four types: semantic ambiguity, visual similarity, imprecise annotations, and boundary targets. First, semantic ambiguity arises when the textual description corresponds to multiple objects within the image, leading the model to segment multiple instances instead of the intended target. Second, visual similarity occurs when the foreground object closely resembles the background in appearance, making it difficult for the model to distinguish between them. Third, imprecise annotations refer to errors introduced by inaccurate or inconsistent labeling, which can mislead the model during training and evaluation. Finally, boundary targets present challenges when the object is located near the image boundary, where limited contextual information may cause the model to misinterpret its position and shape, resulting in inaccurate segmentation.
7 Conclusion
@@
The ability to precisely segment target objects in remote sensing images based on natural language descriptions not only enhances the accessibility of remote sensing data but also opens up new possibilities for interdisciplinary research and practical applications in fields such as disaster response, urban planning, and environmental monitoring. For future work, we aim to embed domain-specific knowledge[42, 11] of remote sensing imagery into language models, such as sensor imaging theory, spatial correlations, and spectral characteristics of ground objects, to further enhance remote sensing data analysis and interpretation. Additionally, another promising research direction is the integration of textual information with remote sensing through geolocation. By incorporating non-traditional geographic data, such as geotagged social media posts[65] and newspaper articles, remote sensing data can be combined with complementary sources. This approach broadens the potential applications[14, 22, 23] of remote sensing visual-language foundation models.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
The authors would like to thank the Editor, Associate Editor, and anonymous reviewers for their helpful comments and suggestions that improved this paper. This work was supported by the Distinguished Young Scholars of Natural Science Foundation of China under Grant 62025107.